OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (6, 1) and has slope 5/6. please show work!

OpenStudy (amistre64):

y = 5/6 x -(5(6)/6)+1 y = 5/6 x -4

OpenStudy (dumbcow):

(y-b) = m(x-a) m=5/6 (a,b) = (6,1) substitute into general formula and get y by itself

OpenStudy (amistre64):

id rather use a formula that already has the algebra worked out :)

OpenStudy (anonymous):

both are right!

OpenStudy (amistre64):

they are both right becasue they are both the same thing; mine just lets you do it all at once

OpenStudy (amistre64):

m(x-a) = mx -ma y-b = mx -ma <=> y = mx -ma +b

OpenStudy (anonymous):

i hate graphs!!

OpenStudy (dumbcow):

yep same formula, his is just already kinda worked out for you so yeah whatever works

OpenStudy (amistre64):

lol.... the screen your looking at right now is a graph :)

OpenStudy (anonymous):

thanks guys!!